# Stochastic gradient descent estimation of generalized matrix factorization models with application to single-cell RNA sequencing data

### v1

@@ Abstract
Single–cell RNA sequencing allows the quantitation of gene expression at the individual cell level, enabling the study of cellular heterogeneity and gene expression dynamics. Dimensionality reduction is a common preprocessing step to simplify the visualization, clustering, and phenotypic characterization of samples. This step, often performed using principal component analysis or closely related methods, is challenging because of the size and complexity of the data. In this work, we present a generalized matrix factorization model assuming a general exponential dispersion family distribution and we show that many of the proposed approaches in the single–cell dimensionality reduction literature can be seen as special cases of this model. Furthermore, we propose a scalable adaptive stochastic gradient descent algorithm that allows us to estimate the model efficiently, enabling the analysis of millions of cells. Our contribution extends to introducing a novel warm start initialization method, designed to accelerate algorithm convergence and increase the precision of final estimates. Moreover, we discuss strategies for dealing with missing values and model selection. We benchmark the proposed algorithm through extensive numerical experiments against state-of-the-art methods and showcase its use in real-world biological applications. The proposed method systematically outperforms existing methods of both generalized and non-negative matrix factorization, demonstrating faster execution times while maintaining, or even enhancing, matrix reconstruction fidelity and accuracy in biological signal extraction. Finally, all the methods discussed here are implemented in an efficient open-source R package, sgdGMF, available at github/CristianCastiglione/sgdGMF.
Keywords: Dimension reduction; Generalized linear models; Matrix factorization; Stochastic optimization; single–cell; RNA–seq.
@@ 1 Introduction
1.1 The ever-increasing single cell RNA sequencing data sets
@@
Single-cell RNA sequencing (scRNA-seq) technologies have revolutionised the comprehension of biological processes by offering a quantitative measure of transcript abundance at the individual cell level. Single-cell resolution is critical for the study of cellular heterogeneity (Wu et al.,, 2023; Kim and Cho,, 2023), temporal dynamics (Kouno et al.,, 2013; Jean-Baptiste et al.,, 2019), and cell-type differentiation (Denyer et al.,, 2019). However, the size and complexity of the data have dramatically increased compared to bulk sample-level assays, challenging statistical methods and software implementations to deal with thousands of genes profiled in millions of cells (Angerer et al.,, 2017; Kharchenko,, 2021).
From a statistical perspective, scRNA-seq yields high-dimensional count data, in which observations (i.e., cells) lie in a $\sim 10{,}000$-dimensional gene space. Working in such high-dimensional spaces poses a wealth of statistical and computational challenges. Hence, one of the first steps in the analysis of single-cell data is dimensionality reduction. Indeed, proper dimensionality reduction reduces both noise and computational complexity and simplifies downstream analyses such as data visualization, cell clustering, and lineage reconstruction (Sun et al.,, 2019; Kiselev et al.,, 2019; Xiang et al.,, 2021). To this end, state-of-the-art methods leverage a low-rank representation of the data to extract useful information while correcting for known confounders such as batch effects or other technical or biological covariates (Risso et al.,, 2018; Townes et al.,, 2019; Agostinis et al.,, 2022).
@@
Yet, the dimensionality reduction of scRNA–seq is not trivial. This technology yields count data with low mean and large variance, leading to extremely skewed distributions with a large fraction of zeros (Hicks et al.,, 2018). To address zero inflation and overdispersion problems, Risso et al., (2018) introduced a zero-inflated negative binomial framework for gene-expression matrix factorization, named ZINB–WaVE. Despite its superior performance compared to other methods, ZINB–WaVE’s computational complexity renders it obsolete for the ever-increasing data volumes (Sun et al.,, 2019; Cao et al.,, 2021). Moreover, in modern UMI-based scRNA–seq datasets the need for modeling zero inflation has decreased (Townes et al.,, 2019; Svensson,, 2020; Nguyen et al.,, 2023; Ahlmann-Eltze and Huber,, 2023). In this regard, simpler models, such as those developed by Townes et al., (2019) and Agostinis et al., (2022), achieve similar performances with a much smaller computational burden. Nevertheless, even these methods struggle to scale with the current massive data volumes.
@@
Consequently, biological researchers often resort to log-transformation of the scRNA–seq count tables and apply conventional principal component analysis (PCA; Ahlmann-Eltze and Huber,, 2023). While this is a much faster alternative, it ignores the complex mean-variance relation and discrete nature of the data and may introduce synthetic biases due to imperfect data transformations (Townes et al.,, 2019). Therefore, the need for fast and memory-efficient matrix factorization tools on the original count scale still remains.
@@ 1.2 A matrix factorization perspective
The dimensionality reduction techniques mentioned above are all examples of matrix factorization, a statistical tool of fundamental importance in many theoretical and applied fields. In general, matrix factorization methods aim at decomposing the target data matrix into the product of two lower-rank matrices explaining the principal modes of variations in the observations. Such low-rank matrices are typically interpreted in terms of factors and loadings. Factors represent stochastic latent variables, i.e. random effects, lying in a small-dimensional space and determining the individual characteristics of each observation in the sample. Loadings are non-stochastic coefficients mapping the latent factors into the observed data space, or some one-to-one transformation thereof. Specifically, in single–cell RNA–seq applications, factors and loadings can be interpreted as “meta genes” and “gene weights”, respectively (Stein-O’Brien et al.,, 2018). In this context, matrix factorization is usually employed to project the cells into the space of the first few factors to then cluster them in discrete groups (Kiselev et al.,, 2019) or infer other low-rank signal, such as pseudotime ordering (Street et al.,, 2018).
@@
Principal component analysis (PCA; Jolliffe,, 1986), i.e., singular value decomposition (SVD), plays a central role in the literature, being the first and most used factorization method proposed in the field. It constitutes the base for several generalizations such as probabilistic PCA (Tipping and Bishop,, 1999), factor analysis, and, more generally, generalized linear latent variable models (GLLVM, Bartholomew et al.,, 2011). Probabilistic formulations equip PCA with a data-generating mechanism that opens the door to alternative inferential procedures, such as likelihood-based and Bayesian approaches. Moreover, it provides a natural way to simulate new synthetic signals using a proper generating mechanism induced by the likelihood specification. Assuming a Gaussian law for the data, PCA can be formulated as the solution to a likelihood maximization problem under appropriate identifiability constraints; see, e.g., Tipping and Bishop, (1999). This perspective unveils why PCA may be suboptimal for non-Gaussian data, such as positive scores, counts, or binary observations.
@@
Over the years, many extensions have been proposed to address the limitations of the Gaussian PCA. Some relevant examples are non-negative matrix factorization (Lee and Seung,, 1999; Wang and Zhang,, 2012), Binary PCA (Schein et al.,, 2003), Poisson PCA (Durif et al.,, 2019; Smallman et al.,, 2020; Kenney et al.,, 2021; Virta and Artemiou,, 2023), exponential family PCA (Collins et al.,, 2001; Mohamed et al.,, 2008; Li and Tao,, 2010; Gopalan et al.,, 2015; Wang et al.,, 2020), generalized linear latent variable models (Niku et al.,, 2017), generalized PCA (Townes et al.,, 2019), generalized matrix factorization (Kidziński et al.,, 2022) and deviance matrix factorization (Wang and Carvalho,, 2023).
@@
Non-negative matrix factorization (NMF; Lee and Seung,, 1999; Wang and Zhang,, 2012) decomposes positive score matrices by minimizing either the squared error loss or the Kullbak-Leibler loss under non-negativity constraints for the factor and loading matrices. The exploitation of non-negative patterns proved successful in many applied fields, such as computer vision and recommendation systems, as well as omics feature extraction (Stein-O’Brien et al.,, 2018).
@@ 1.3 Our contribution
Table 1: List of the major matrix factorization models (first column) available in the literature for exponential family data. For each model, we report the corresponding R package (second column) and we describe its characteristics (from the third to the last columns). Each feature is marked with ✓ if it is completely implemented in the package and $\times$ otherwise. The column Families lists the most common distributions belonging to the exponential family along with some generalizations: Normal (N), Gamma (G), Binomial (B), Poisson (P), Negative Binomial (NB), quasi-likelihood (Q) and zero-inflated (ZI) models. The column Effects refers to the regression effects that can be included in the linear predictor and uses the notation introduced in Section 2, Equation (3). The column Implementation describes some technical features of the numerical implementation, such as the language used for the core computations (Core), if parallel computing is allowed (Parallel) and if minibatch subsample and stochastic optimization methods are available (Stochastic).
@@
N G B P NB Q ZI $\mathbf{X}\mathbf{B}^{\scriptscriptstyle\top}$ $\boldsymbol{\Gamma}\mathbf{Z}^{\scriptscriptstyle\top}$ $\mathbf{U}\mathbf{V}^{\scriptscriptstyle\top}$ Core Parallel Stochastic
RSpectra PCA $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ F/C++ $\times$
glmpca GMF $\times$ $\times$ $\times$ $\times$ R $\times$
zinbwave GMF $\times$ $\times$ $\times$ $\times$ $\times$ R $\times$
newwave GMF $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ R
@@
dmf GMF $\times$ $\times$ R $\times$ $\times$
@@
gllvm GLLVM $\times$ $\times$ R $\times$ $\times$
NMF NMF $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ C++ $\times$
NNLM NMF $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ C++ $\times$
cmfrec NMF $\times$ $\times$ $\times$ $\times$ $\times$ C++ $\times$
Alongside, we propose an efficient R/C++ implementation of the proposed method in the new open-source R package sgdGMF, freely available on GitHub [^1] Refer to the GitHub repository github/CristianCastiglione/sgdGMF. Compared to alternative implementations in R, sgdGMF offers one of the most complete and flexible estimation frameworks for generalized matrix factorization modelling, allowing for all standard exponential family distributions, quasi-likelihood models, row- and column-specific regression effects, as well as model-based missing value imputation. Moreover, it provides several algorithms for parameter estimation, including the proposed stochastic gradient descent approach. To enhance scalability, parallel computing is employed as much as possible at any stage of the analysis, including initialization, estimation, model selection and post-processing. Table 1 compares all these features with alternative packages freely available in R.
We showcase the sgdGMF implementation on both simulated and real data, demonstrating the scalability of the proposed method on gene expression matrices of different dimensions. In all the numerical experiments we consider, the proposed stochastic gradient approach outperforms the alternative state-of-the-art methods in terms of execution time while having superior signal reconstruction quality, measured as out-of-sample residual deviance and cluster separation in the latent space.
The paper is organized as follows. In Section 2 we formally define the class of generalized matrix factorization models, we formulate the associated estimation problem and we discuss the connections with other models in the literature. In Section 3, we introduce the proposed stochastic optimization method building upon quasi-Newton and stochastic gradient descent algorithms; additionally, we propose an efficient approach for parameter initialization. In Section 4, we discuss different methods for model selection. In Section 5, we empirically compare the proposed algorithm with several state-of-the-art methods in the literature through an extensive simulation study. In Section 6, we propose two case studies on real datasets of medium- and high-dimensional sizes, respectively, proving the effectiveness of our approach to extract coherent biological signals while maintaining a high level of computational efficiency. Section 7 is devoted to a concluding discussion and future research directions.
@@ 2 Model specification
where $b(\cdot)$ and $c(\cdot,\cdot)$ are family-specific functions. Specifically, $b(\cdot)$ is a convex twice differentiable cumulant function and $c(\cdot,\cdot)$ is the so-called log-partition function. By the fundamental properties of the dispersion exponential family, the mean and variance of $y_{ij}$ satisfy the identities $\mu_{ij}=\accentset{\mbox{\large.}}{b}(\theta_{ij})$ and $\nu(\mu_{ij})=\accentset{\mbox{\large..}}{b}(\theta_{ij})$, which imply $\nu(\cdot)=(\accentset{\mbox{\large..}}{b}\circ\accentset{\mbox{\large.}}{b}^{-1})(\cdot)$, where $\accentset{\mbox{\large.}}{b}(\cdot)$ and $\accentset{\mbox{\large..}}{b}(\cdot)$ denote the first and second derivatives of $b(\cdot)$, respectively. Therefore, $\accentset{\mbox{\large.}}{b}(\cdot)$ is a bijective map and $\theta_{ij}$ is uniquely determined by $\mu_{ij}$
Some relevant distributions belonging to the exponential family are, among others, the Gaussian, Inverse Gaussian, Gamma, Poisson, Binomial and Negative Binomial laws. Table 2 shows the family-specific variance functions related to these examples as well as the associated canonical link and deviance functions. We recall that the deviance function of model (2) is defined as $D(y,\mu)=-2\log\{f_{\phi}(y,\mu)/f_{\phi}(y,y)\}$, where $f_{\phi}(y,\mu)=f(y;\theta,\phi)$ denotes the likelihood function relative to observation $y$ expressed as a function of mean $\mu$ and possibly depending on dispersion $\phi$.
Table 2: Exponential family laws along with their support space, canonical link ($g_{c}$), variance ($\nu$), and rescaled deviance ($D/2a$) functions. Here, we denote by $N\in\mathbb{N}$ and $\alpha>0$ the size and variance parameters of the Binomial and Negative Binomial distributions, respectively.
Distribution Support $g_{c}(\mu)$ $\nu(\mu)$ $D(y,\mu)/2a(\phi)$
Gaussian $\mathbb{R}$ $\mu$ $1$ $(y-\mu)^{2}/2$
Gamma $\mathbb{R}_{+}$ $1/\mu$ $\mu^{2}$ $(y-\mu)/\mu-\log(y/\mu)$
Inv. Gauss. $\mathbb{R}_{+}$ $1/\mu^{2}$ $\mu^{3}$ $(y-\mu)^{2}/(2y\mu^{2})$
Poisson $\mathbb{N}_{+}$ $\log(\mu)$ $\mu$ $y\log(y/\mu)-(y-\mu)$
@@
where $g(\cdot)$ is a continuously differentiable bijective link function and $\eta_{ij}$ is a linear predictor. The latter is represented as an additive decomposition of three terms: a column-specific regression effect, $(\mathbf{X}\mathbf{B}^{\top})_{ij}=\mathbf{x}_{i:}^{\top}\boldsymbol{\beta}_{:j}$, a row-specific regression effect, $(\boldsymbol{\Gamma}\mathbf{Z}^{\top})_{ij}=\boldsymbol{\gamma}_{i:}^{\top}\mathbf{z}_{:j}$, and a residual matrix factorization $(\mathbf{U}\mathbf{V}^{\top})_{ij}=\mathbf{u}_{i:}^{\top}\mathbf{v}_{:j}$. Specifically, $\mathbf{x}_{i:}\in\mathbb{R}^{p}$ and $\mathbf{z}_{:j}\in\mathbb{R}^{q}$ denote observed covariate vectors, $\boldsymbol{\beta}_{:j}\in\mathbb{R}^{p}$ and $\boldsymbol{\gamma}_{i:}\in\mathbb{R}^{q}$ are unknown regression coefficient vectors, while $\mathbf{u}_{i:}\in\mathbb{R}^{d}$ and $\mathbf{v}_{:j}\in\mathbb{R}^{d}$ encode latent traits explaining the residual modes of variation in the data that are not captured by the regression effects. Notice that the linear predictor can be equivalently expressed using the matrix decomposition $\eta_{ij}=(\mathbf{U}_{\star}\mathbf{V}_{\star}^{\top})_{ij}$, where $\mathbf{U}_{\star}=[\mathbf{X},\boldsymbol{\Gamma},\mathbf{U}]$ and $\mathbf{V}_{\star}=[\mathbf{B},\mathbf{Z},\mathbf{V}]$ are partially observed low-rank matrices. In the following sections, and in particular in Section 3, we make extensive use of this more compact form hiding the star-subscript to lighten the notation and simplify the exposition. Finally, we introduce the vector of unknown parameters in the model as $\boldsymbol{\psi}=(\boldsymbol{\beta}^{\top},\boldsymbol{\gamma}^{\top},\mathbf{u}^{\top},\mathbf{v}^{\top},\phi)^{\top}$, where the lower-case letters represent the flat vectorization of the corresponding matrix forms; for instance, $\boldsymbol{\beta}=\textrm{{vec}}(\mathbf{B})$.

### v2

@@ Abstract
Single-cell RNA sequencing allows the quantification of gene expression at the individual cell level, enabling the study of cellular heterogeneity and gene expression dynamics. Dimensionality reduction is a common preprocessing step critical for the visualization, clustering, and phenotypic characterization of samples. This step, often performed using principal component analysis or closely related methods, is challenging because of the size and complexity of the data. In this work, we present a generalized matrix factorization model assuming a general exponential dispersion family distribution and we show that many of the proposed approaches in the single-cell dimensionality reduction literature can be seen as special cases of this model. Furthermore, we propose a scalable adaptive stochastic gradient descent algorithm that allows us to estimate the model efficiently, enabling the analysis of millions of cells. We benchmark the proposed algorithm through extensive numerical experiments against state-of-the-art methods and showcase its use in real-world biological applications. The proposed method systematically outperforms existing methods of both generalized and non-negative matrix factorization, demonstrating faster execution times and parsimonious memory usage, while maintaining, or even enhancing, matrix reconstruction fidelity and accuracy in biological signal extraction. On real data, we show that our method scales seamlessly to millions of cells, enabling dimensionality reduction in large single-cell datasets. Finally, all the methods discussed here are implemented in an efficient open-source R package, sgdGMF, available on CRAN.
Keywords: Dimension reduction; Generalized linear models; Matrix factorization; Stochastic optimization; single-cell; RNA-seq.
@@ 1 Introduction
1.1 The ever-increasing single-cell RNA sequencing data sets
@@
Single-cell RNA sequencing (scRNA-seq) technologies have revolutionized the comprehension of biological processes by offering a quantitative measure of transcript abundance at the individual cell level. Single-cell resolution is critical for the study of cellular heterogeneity (Wu et al.,, 2023; Kim and Cho,, 2023), temporal dynamics (Kouno et al.,, 2013; Jean-Baptiste et al.,, 2019), and cell-type differentiation (Denyer et al.,, 2019). However, the size and complexity of the data have dramatically increased compared to bulk sample-level assays, challenging statistical methods and software implementations to deal with thousands of genes profiled in millions of cells (Angerer et al.,, 2017; Kharchenko,, 2021).
From a statistical perspective, scRNA-seq yields high-dimensional count data, in which observations (i.e., cells) lie in a $\sim 10{,}000$-dimensional gene space. Working in such high-dimensional spaces poses a wealth of statistical and computational challenges. Lähnemann et al., (2020) identified eleven grand challenges in single-cell data science; here we focus on three of them: (i) the handling of data sparsity, (ii) the integration of datasets across samples and experiments, and (iii) the definition of flexible statistical frameworks for the discovery of complex gene expression patterns. These challenges are exacerbated by the growing size and complexity of single-cell datasets. Indeed, while early studies focused on few cells from one or few samples, modern studies involve complex experimental designs, collecting thousands of cells from several individuals across different experimental conditions (see, e.g., Stephenson et al.,, 2021; Perez et al.,, 2022).
@@
Regarding sparsity, single-cell sequencing technologies yield count data with low mean and large variance, leading to extremely skewed distributions with a large fraction of zeros (Hicks et al.,, 2018). To address zero inflation and overdispersion problems, Risso et al., (2018) introduced a zero-inflated negative binomial framework for gene-expression matrix factorization, named ZINB-WaVE. Despite its superior performance compared to other methods, ZINB-WaVE’s computational complexity renders it obsolete for the ever-increasing data volumes (Sun et al.,, 2019; Cao et al.,, 2021). Moreover, in modern UMI-based scRNA-seq datasets, the need for modeling zero inflation has decreased (Townes et al.,, 2019; Svensson,, 2020; Nguyen et al.,, 2023; Ahlmann-Eltze and Huber,, 2023). In this regard, simpler Negative Binomial models, such as those developed by Townes et al., (2019) and Agostinis et al., (2022), achieve similar performances with a much smaller computational burden. Nevertheless, even these methods struggle to scale with the current massive data volumes.
Another main challenge is the integration of data across different samples, experiments, or labs. These effects, often globally referred to as batches, represent a source of unwanted variation that need to be accounted for in the analysis. In unsupervised settings, for instance cell clustering or trajectory analysis, batch effects are usually corrected for with ad hoc methods (see, e.g., Haghverdi et al.,, 2018; Korsunsky et al.,, 2019). While many such methods exist, the batch integration of single-cell data is still an open problem (Luecken et al.,, 2025; Lähnemann et al.,, 2020).
Factor analysis models are a promising, unified framework that allows to account for the count nature of the data, adjust for potential sources of unwanted variation, and provide a parsimonious, low-dimensional representation that can be used as input for downstream analyses. To this end, state-of-the-art methods leverage a low-rank representation of the data to extract useful information while correcting for known confounders such as batch effects or other technical or biological covariates (Risso et al.,, 2018; Townes et al.,, 2019; Agostinis et al.,, 2022).
@@
However, because of the lack of scalability of state-of-the-art count-based factor models, biological researchers often resort to log-transformation of the scRNA-seq count tables and apply conventional principal component analysis (PCA; Ahlmann-Eltze and Huber,, 2023). While this is a much faster alternative, it ignores the complex mean-variance relation and discrete nature of the data and may introduce synthetic biases due to imperfect data transformations (Townes et al.,, 2019). Furthermore, simple PCA of log-transformed counts fails to account for batch effects. Therefore, the need still remains for fast and memory-efficient matrix factorization tools on the original count scale able to include covariates, such as batch effects.
@@ 1.2 A matrix factorization perspective
The above-mentioned factor analysis models are examples of matrix factorization, a statistical tool of fundamental importance in many theoretical and applied fields. In general, matrix factorization methods aim at decomposing the target data matrix into the product of two lower-rank matrices explaining the principal modes of variation in the observations. Such low-rank matrices are typically interpreted in terms of factors and loadings. Factors represent stochastic latent variables, i.e., random effects, lying in a small-dimensional space and determining the individual characteristics of each observation in the sample. Loadings are non-stochastic coefficients mapping the latent factors into the observed data space, or some one-to-one transformation thereof. Specifically, in single-cell RNA-seq applications, factors and loadings can be interpreted as meta-genes and gene-weights, respectively (Brunet et al.,, 2004; Stein-O’Brien et al.,, 2018). In this context, matrix factorization is usually employed to project the cells into the space of the first few factors to then cluster them in discrete groups (Kiselev et al.,, 2019) or infer other low-rank signal, such as pseudotime ordering (Street et al.,, 2018).
@@
Principal component analysis (PCA; Jolliffe,, 1986), i.e., singular value decomposition (SVD), plays a central role in the literature, being the first and most used factorization method proposed in the field. It constitutes the base for several generalizations such as probabilistic PCA (Tipping and Bishop,, 1999), factor analysis, and, more generally, generalized linear latent variable models (GLLVM, Bartholomew et al.,, 2011). Probabilistic formulations equip PCA with a data-generating mechanism that opens the door to alternative inferential procedures, such as likelihood-based and Bayesian approaches. Moreover, it provides a natural way to simulate new synthetic signals using a proper generating mechanism induced by the likelihood specification. Assuming a Gaussian law for the data, PCA can be formulated as the solution to a likelihood maximization problem under appropriate identifiability constraints (see, e.g., Tipping and Bishop,, 1999). This perspective unveils why PCA may be suboptimal for non-Gaussian data, such as positive scores, counts, or binary observations.
@@
Over the years, many extensions have been proposed to address the limitations of the Gaussian PCA. Some relevant examples are non-negative matrix factorization (Lee and Seung,, 1999; Wang and Zhang,, 2012), Binary PCA (Schein et al.,, 2003), Poisson PCA (Durif et al.,, 2019; Smallman et al.,, 2020; Kenney et al.,, 2021; Virta and Artemiou,, 2023), exponential family PCA (Collins et al.,, 2001; Mohamed et al.,, 2008; Li and Tao,, 2010; Gopalan et al.,, 2015; Wang et al.,, 2020), generalized linear latent variable models (Niku et al.,, 2017), generalized factor model (Liu et al.,, 2023; Nie et al.,, 2024), covariate-augmented overdispersed Poisson factor model (Liu and Zhong,, 2024), generalized PCA (Townes et al.,, 2019), generalized matrix factorization (Kidziński et al.,, 2022) and deviance matrix factorization (Wang and Carvalho,, 2023).
@@
Non-negative matrix factorization (NMF; Lee and Seung,, 1999; Wang and Zhang,, 2012) decomposes positive score matrices by minimizing either the squared error loss or the Kullbak–Leibler loss under non-negativity constraints for the factor and loading matrices. The exploitation of non-negative patterns proved successful in many applied fields, such as computer vision and recommendation systems, as well as omics feature extraction (Brunet et al.,, 2004; Stein-O’Brien et al.,, 2018).
@@ 1.3 Our contribution
Table 1: List of the major matrix factorization models available in the literature for exponential family data. For each model, we report the corresponding R package (first column) and we describe its characteristics. Each feature is marked with ✓ if it is completely implemented in the package and $\times$ otherwise. The column Model indicates the broad model family implemented in the package, where GMF and NMF stand for generalized matrix factorization, and non-negative matrix factorization, respectively. The column Families lists the most common distributions belonging to the exponential family along with some generalizations: Normal (N), Gamma (G), Binomial (B), Poisson (P), Negative Binomial (NB), quasi-likelihood (Q) and zero-inflated (ZI) models. The column Effects refers to the regression effects that can be included in the linear predictor and uses the notation introduced in Section 2, Equation (3). The column Implementation describes some technical features of the numerical implementation, such as the language used for the core computations (Core), if automatic missing value estimation is allowed (Missing), if parallel computing is allowed (Parallel), and if minibatch subsample and stochastic optimization methods are available (Stochastic).
@@
N G B P NB Q ZI $\mathbf{X}\mathbf{B}^{\scriptscriptstyle\top}$ $\boldsymbol{\Gamma}\mathbf{Z}^{\scriptscriptstyle\top}$ $\mathbf{U}\mathbf{V}^{\scriptscriptstyle\top}$ Core Missing Parallel Stochastic
RSpectra PCA $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ F/C++ $\times$ $\times$
NMF NMF $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ C++ $\times$ $\times$
NNLM NMF $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ C++ $\times$
cmfrec NMF $\times$ $\times$ $\times$ $\times$ $\times$ C++ $\times$ $\times$
gllvm GMF $\times$ $\times$ R $\times$ $\times$ $\times$
GFM GMF $\times$ $\times$ $\times$ $\times$ $\times$ C++ $\times$ $\times$ $\times$
COAP GMF $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ C++ $\times$ $\times$ $\times$
glmpca GMF $\times$ $\times$ $\times$ $\times$ R $\times$ $\times$
zinbwave GMF $\times$ $\times$ $\times$ $\times$ $\times$ R $\times$ $\times$
newwave GMF $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ R $\times$
@@
dmf GMF $\times$ $\times$ R $\times$ $\times$ $\times$
@@
Alongside, we propose an efficient R/C++ implementation of the proposed method in the new open-source R package sgdGMF, freely available on CRAN [^1] See https://CRAN.R-project.org/package=sgdGMF. Compared to alternative implementations in R, sgdGMF offers one of the most complete and flexible estimation frameworks for generalized matrix factorization modelling, allowing for all standard exponential family distributions, quasi-likelihood models, row- and column-specific regression effects, as well as model-based missing value imputation. Moreover, it provides several algorithms for parameter estimation, including the proposed stochastic gradient descent approach. To enhance scalability, parallel computing is employed as much as possible at any stage of the analysis, including initialization, estimation, model selection and post-processing. Table 1 compares all these features with alternative packages freely available in R.
We showcase the sgdGMF implementation on both simulated and real data, demonstrating the scalability of the proposed method on gene expression matrices of different dimensions. In all the numerical experiments we consider, the proposed stochastic gradient approach outperforms the alternative state-of-the-art methods in terms of execution time while having superior signal reconstruction quality, measured as out-of-sample residual deviance, logarithmic root mean squared error, and cluster separation in the latent space.
The paper is organized as follows. In Section 2 we formally define the class of generalized matrix factorization models, we formulate the associated estimation problem and we discuss the connections with other models in the literature. In Section 3, we introduce the proposed stochastic optimization method building upon quasi-Newton and stochastic gradient descent algorithms; additionally, we propose an efficient approach for parameter initialization. In Section 4, we briefly discuss some additional computational aspects, such as initialization and model selection. In Section 5, we empirically compare the proposed algorithm with several state-of-the-art methods in the literature through an extensive simulation study. In Section 6, we propose two case studies on real datasets of medium- and high-dimensional sizes, respectively, proving the effectiveness of our approach to extract coherent biological signals while maintaining a high level of computational efficiency. Section 7 is devoted to a concluding discussion and future research directions.
@@ 2 Model specification
[…]
Some relevant distributions belonging to the exponential family are, among others, the Gaussian, Inverse Gaussian, Gamma, Poisson, Binomial, and Negative Binomial laws (see Supplementary Table 1 for family-specific variance functions and the associated canonical link and deviance functions of these examples). Of particular interest for omics applications are the Poisson and Negative Binomial distributions, as the gene expressions are typically obtained from technologies that yield count data (e.g., RNA sequencing or in-situ hybridization). However, formulating the model in the general form of the exponential family allows for a more general treatment and for the application to other omics technologies, such as mass spectrometry and microarray assays, which yield continuous readouts.
@@
where $g(\cdot)$ is a continuously differentiable bijective link function and $\eta_{ij}$ is a linear predictor. The latter is represented as an additive decomposition of three terms: a column-specific regression effect, $(\mathbf{X}\mathbf{B}^{\top})_{ij}=\mathbf{x}_{i:}^{\top}\boldsymbol{\beta}_{:j}$, a row-specific regression effect, $(\boldsymbol{\Gamma}\mathbf{Z}^{\top})_{ij}=\boldsymbol{\gamma}_{i:}^{\top}\mathbf{z}_{:j}$, and a residual matrix factorization $(\mathbf{U}\mathbf{V}^{\top})_{ij}=\mathbf{u}_{i:}^{\top}\mathbf{v}_{:j}$. Specifically, $\mathbf{x}_{i:}\in\mathbb{R}^{p}$ and $\mathbf{z}_{:j}\in\mathbb{R}^{q}$ denote observed covariate vectors, $\boldsymbol{\beta}_{:j}\in\mathbb{R}^{p}$ and $\boldsymbol{\gamma}_{i:}\in\mathbb{R}^{q}$ are unknown regression coefficient vectors, while $\mathbf{u}_{i:}\in\mathbb{R}^{d}$ and $\mathbf{v}_{:j}\in\mathbb{R}^{d}$ encode latent traits explaining the residual modes of variation in the data that are not captured by the regression effects. Finally, we introduce the vector of unknown parameters in the model as $\boldsymbol{\psi}=(\boldsymbol{\beta}^{\top},\boldsymbol{\gamma}^{\top},\mathbf{u}^{\top},\mathbf{v}^{\top},\phi)^{\top}$, where the lower-case letters represent the flat vectorization of the corresponding matrix forms; for instance, $\boldsymbol{\beta}=\textrm{{vec}}(\mathbf{B})\in\mathbb{R}^{pm}$.
Specifically, in the scRNA-seq context, the regression term $\mathbf{x}_{i:}^{\top}\boldsymbol{\beta}_{:j}$ is often used to control for cell-specific technical confounders, such as batch effects and individual characteristics in multi-subject studies, with $\mathbf{x}_{i:}$ containing indicator variables of the group assignment and group-specific attributes. Similarly, the second regression term, $\boldsymbol{\gamma}_{i:}^{\top}\mathbf{z}_{:j}$, can account for gene-specific information contained in the covariate vector $\mathbf{z}_{:j}$, such as GC-content or known functional interactions between genes. Notice that, even in the absence of external confounders, such regression effects allow for the inclusion of cell- and gene-specific intercepts, say $\eta_{ij}=\beta_{0j}+\gamma_{0i}+\mathbf{u}_{i:}^{\top}\mathbf{v}_{j:}$, which play the role of centering factors in the linear predictor metric. In scRNA-seq, the cell-specific intercept $\gamma_{0i}$ is of particular importance, as it acts as a scale factor that allows the model to be applied to raw counts rather than library size normalized data (Risso et al.,, 2018), which is critical since normalization can lead to distortions (Townes et al.,, 2019).
The latent traits $\mathbf{u}_{i:}$ can be interpreted as meta-gene variables representing the fundamental biological characteristics of the $i$th cell in a latent low-dimensional space that accounts for the covariates effects, with interpretation similar to usual PCA. For example, the meta-gene variables can be used for data visualization, cell clustering, and lineage reconstruction. It is worth noting that, unlike formulation (3), most factorization models proposed in the literature and implemented in standard computing environments do not support the inclusion of row effects $\mathbf{x}_{i:}^{\top}\boldsymbol{\beta}_{:j}$’s, column effects $\boldsymbol{\gamma}_{i:}^{\top}\mathbf{z}_{:j}$’s, or both, resulting in latent representations that still capture variability of unwanted confounders, such as batch effects, or that do not include a normalization factor. This limitation applies to gmf (Kidziński et al.,, 2022), dmf (Wang and Carvalho,, 2023), GFM (Liu et al.,, 2023; Nie et al.,, 2024), COAP (Liu and Zhong,, 2024), NMF (Gaujoux and Seoighe,, 2010), and NNLM (Lin and Boutros,, 2020), as summarized in Table 1.
Following the naming convention introduced by Kidziński et al., (2022), we refer to the model specification presented so far in (1)–(3) as generalized matrix factorization (GMF). Alternative nomenclatures, such as exponential family principal component analysis (EPCA; Collins et al.,, 2001; Mohamed et al.,, 2008; Li and Tao,, 2010), generalized low-rank models (GLRM; Udell et al.,, 2016), generalized linear latent variable model (GLLVM; Niku et al.,, 2017; Hui et al.,, 2017), generalized principal component analysis (glmPCA; Townes et al.,, 2019), deviance matrix factorization (DMF; Wang and Carvalho,, 2023), or generalized factor model (GFM; Liu et al.,, 2023; Liu and Zhong,, 2024; Nie et al.,, 2024), can also be found in the literature. Straightforward extensions of the GMF specification include pseudo-likelihood models, with $f(y_{ij};\boldsymbol{\psi})=\exp\{-L_{\phi}(y_{ij},\eta_{ij})\}$ being the negative exponential of a loss function $L_{\phi}(\cdot,\cdot):\mathcal{Y}\times\mathbb{R}\rightarrow\mathbb{R}_{+}$. Also, vector generalized estimating equations for overdispersed data are a particular case of this more general setup.
2.1 Parameter identifiability